Eva is 6 years older than her brother. In 4 years, Eva will be 26 more than the half of her brother's age. How old are they now?
Answer by mananth(16946)

brother's age x

brother's age x
Eva x+6
4 years later
brother's age x+4
Eva x+6+4=x+10
x+10=1/2(x+4)+26
multiply by 2
2(x+10)=x+4 +52
2x+20=x+56
2x-x=56-20
2x=36
/2
x=18 brother's age
Eva= 18+6 = 24years